OpenStudy (anonymous):

If the volume of a cone is 12pie cubic inches when the radius is 2 inches and the height is 9inches, find the volume of a cone when then radius is 3inches and height is 6inches.

OpenStudy (anonymous):

vol. of cone=(1/3)pi*(r^2)*h symbols having their usual meaning. just substitute the values in the formula and get the answer 18pi

OpenStudy (anonymous):

v=(1/3)pi*(3^2)*6=18pi cubic inches